# Do Retrieval-Augmented Language Models Adapt to Varying User Needs?

### v1

@@ Abstract
Recent advancements in Retrieval-Augmented Language Models (RALMs) have demonstrated their efficacy in knowledge-intensive tasks. However, existing evaluation benchmarks often assume a single optimal approach to leveraging retrieved information, failing to account for varying user needs. This paper introduces a novel evaluation framework that systematically assesses RALMs under three user need cases—Context-Exclusive, Context-First, and Memory-First—across three distinct context settings: Context Matching, Knowledge Conflict, and Information Irrelevant. By varying both user instructions and the nature of retrieved information, our approach captures the complexities of real-world applications where models must adapt to diverse user requirements. Through extensive experiments on multiple QA datasets, including HotpotQA, DisentQA, and our newly constructed synthetic URAQ dataset, we find that restricting memory usage improves robustness in adversarial retrieval conditions but decreases peak performance with ideal retrieval results and model family dominates behavioral differences. Our findings highlight the necessity of user-centric evaluations in the development of retrieval-augmented systems and provide insights into optimizing model performance across varied retrieval contexts. We will release our code and URAQ dataset upon acceptance of the paper.
@@ 1 Introduction
Recent advances in Language Models (LMs) have yielded impressive performance in knowledge-intensive tasks through Retrieval Augmented Generation (RAG) Lewis et al. (2020), including Real-time Question Answering Wang et al. (2024b), Educational Tutoring Han et al. (2024), and Personal Assistants Wang et al. (2024c). While these applications showcase RAG’s versatility, they also demand LMs that can adapt to diverse user needs—expressed via instructions on whether to prioritize external evidence or internal knowledge. For instance, Real-time QA may rely heavily on updated external facts, whereas tutoring may draw more on the model’s conceptual understanding. Despite this potential, current RAG methods still struggle with identifying relevant references Laban et al. (2024), resolving knowledge conflicts Wang et al. (2024a), and reasoning effectively Islam et al. (2024). These challenges underscore the need for robust evaluation strategies capturing how well Retrieval Augmented Language Models (RALMs) adapt to evolving user requirements.
@@
Even though existing RAG/RALM benchmarks (Yu et al., 2024; Es et al., 2023; Chen et al., 2024)—including those that focus on multi-scenario evaluations (Friel et al., 2024; Zhu et al., 2024)—have advanced retrieval-augmented evaluation, they typically assume a single “optimal” approach to external information (e.g., always relying on retrieved context). This narrow perspective overlooks how diverse user instructions can dramatically alter model behavior and performance within the same scenario. In medical fact-checking, for instance, one user might demand answers derived only from peer-reviewed studies, while another relies on the model’s internal knowledge—even if these sources conflict (Miao et al., 2024). Such constraints underscore an urgent question: how can we systematically evaluate LMs under varying context usage requirements to reflect different user needs?
In this paper, we present a simple yet effective evaluation framework that rigorously examines how Retrieval-Augmented Language Models (RALMs) respond to varying user instructions and context conditions. We consider three generic user cases(1) Context-Exclusive , (2) Context-First, and (3) Memory-First —to capture different degrees of reliance on external information versus internal knowledge. Alongside these cases, we vary the context settings(a) Context Matching, (b) Knowledge Conflict, and (c) Information Irrelevant—to represent scenarios where retrieved materials may align with, contradict, or fail to address the query. By intersecting user cases with distinct context conditions, we more closely mirror the complexities of real-world applications, where both the user’s priorities and the reliability of retrieved information can shift dramatically. This approach reveals how each scenario might alter the correct response—especially when context and memory conflict—an aspect often overlooked in previous work.
@@ 2.3 Recent RAG Benchmark
Previous RAG benchmarks like RAGAS Es et al. (2023) and RGB Chen et al. (2024) have facilitated progress by quantifying performance across various scenarios. However, many of these benchmarks focused on a single type of optimal setting in terms of context usages (for instance, always prioritizing the context), overlooking how different user instructions may drastically affect model behaviors and performances. Moreover, previous multi-scenario evaluations (Friel et al. 2024; Zhu et al. 2024), while covering a wide range of specific tasks and purpose abundant metrics for evaluating different aspects of RAG systems, also tend to follow the paradigm of focusing on singular optimality, neglecting that different user needs can actually happen in the same scenario, ultimately hindering the comprehensiveness of benchmark. Our work diverges by decoupling evaluation criteria from predefined singular optimality and measuring model capability to adapt to dynamic user needs. This mirrors real-world deployments where systems must honor diverse users’ requirements rather than optimize for monolithic accuracy.
@@ 3.1 User Need Cases
Figure 2: An illustration of the framework with an example question with its possible retrieved context and the ground truth answer under each situation. According to different user needs and context settings, the ground truth answer can be different.
@@ 3.2 Context Settings
All retrieved contexts are unrelated to the question. This setting simulates the Needle-In-a-Haystack Laban et al. (2024) situation and tests the model’s ability on knowledge selection. Models are expected to avoid hallucinating and admit knowledge gaps by responding with “I don’t know” under Case 1 instructions or relying on its memory in Case 2 and 3.
@@ 4.1 Datasets
Overview of QA Datasets.
This experiment employs three QA datasets: HotpotQA Yang et al. (2018), DisentQA Neeman et al. (2023), and our synthetic User-focused Retrieval-Augmented QA (URAQ). To assess RALMs’ real-world performance, we use HotpotQA and DisentQA versions augmented with conflicted knowledge by Shaier et al. (2024) for the retrieval-content knowledge conflict setting. While valuable, these benchmarks lack controlled knowledge boundaries and have varying question difficulty, limiting evaluation. They also rely on long-document contexts, restricting retrieval diversity. URAQ addresses these issues with uniformly difficult questions and concise factual contexts, enabling evaluation under extensive retrieval without exceeding LMs’ context windows.
@@
Table 1: Basic information of the three datasets used in the experiment. For the sequence of the number of retrieved context,the number of retrieved context is increased in a exponential way until the average number of tokens at the highest number of each sequence reaches around 20k in order to balance the effectiveness of the experiment on long context and the consumption of computational resources. The Max. Token, which refers to the number of maximum tokens among all samples for a dataset, may vary based on context retrieved.
URAQ Construction.
We construct URAQ by first generating simple, distinct knowledge statements via GPT-4o-mini OpenAI et al. (2024) and removing near-duplicates using SentenceBERT Reimers and Gurevych (2019), then creating both original and “manipulated” versions by substituting key information or adding negations. For each knowledge pair, we produce a question requiring 1–5 reasoning steps and two separate answers (one from the original knowledge, one from the manipulated), ultimately selecting the 4-hop subset for the final dataset. A detailed description of this procedure is provided in Appendix A, ensuring the pipeline’s applicability across various domains.
4.2 Context Setting and Prompt Formatting
Retrieval Context Setting.
To examine how performance changes with varying amounts of retrieved context, rather than using a fixed retrieval count as in previous work Zhu et al. (2024), we evaluate LM performance by exponentially increasing the retrieval count across different datasets, shown in Table 1. To assess the models’ tolerance to distracting or irrelevant contexts, we ensure that only one relevant context is present for both the context-matching and conflicting settings, randomly positioned within the prompt. All other contexts are selected from a pool of original and manipulated knowledge that excludes any information directly related to the current question.
Prompt Formatting
[…]
where <sys></sys> and <user></user> denote the system prompt and the user prompt. Among all data samples, the $I_{u}$ and $C$ may change according to the user case and context setting, while the $I_{f}$ remaining the same by instructing models to directly output a simple answer that is either a numeric value, a boolean ("yes" or "no"), or an entity, as described in Section 4. A complete example prompt template is in the Appendix C.
@@ 4.3 Evaluation Metrics
To rigorously assess user-need awareness across across different user needs with different retrieval content, we test each user need with identical questions but varying the guidance on context usage, spanning three levels:
@@ 5.2 General Performance for Each User Need
(1) Case-Level Accuracy curve of Qwen2.5 on HotpotQA dataset.
(2) Case-Level Accuracy curve of Llama-3.1 on HotpotQA dataset.
Figure 4: Case-Level Accuracy curve of Qwen2.5 and Llama-3.1-70B on HotpotQA
@@ 5.3.1 Performance on Optimal Context
Table 2: Percentage of errors that is "I don’t know" among the shortest 100 randomly selected samples that under Context Matching setting that is incorrect for Context-Exclusive user need and correct for Context or Memory-First. A number exceeding 50 hints that the model is leaning towards reject answering when it has trouble locating the source or deducing the answers.
@@ 6 Conclusion
We introduce an evaluation framework for RALMs that systematically assesses performance across diverse user needs and context settings. By decomposing user instructions into three generic user need cases (Context-only, Context-priority, Memory-priority) and three context settings (Match, Conflict, Irrelevant), our framework provides comprehensive insights into model capabilities and limitations. Our analysis covers overall user requirements, case-level evaluations, and the impact of varying context contents across different context lengths. The findings highlight the need for user-centric evaluations and architectural innovations to enhance RAG system reliability and real-world applicability.
@@ B.3 Memory-First
Appendix C Example Input Prompt
In this section, we introduce an example input prompt that is designed for Case 1 Setting a with 2 total retrieved context following the abstract input $(I_{f},I_{u},C,Q)$ in Section 4.2. The prompt is formatted with XML for both input and output. Specifically, the formatting instructions $I_{f}$ are separated into two parts: 1) The first and second instructions in the system prompt describing that the answer should be as simple as possible with XML format. 2) The instruction in the user prompt about format of context with an reinforcement of output format. The user need instruction $I_{u}$ is at the third instruction in the system prompt. The retrieved context $C$ is all the sentences in the user prompt within <info> xml tag. Finally, the question $Q$ is at the last of user prompt within <question> xml tag.

### v2

@@ Abstract
Recent advancements in Retrieval-Augmented Language Models (RALMs) have demonstrated their efficacy in knowledge-intensive tasks. However, existing benchmarks often assume a singular view of optimal information use, neglecting diverse user needs where ’correctness’ can mean faithfulness to instructed sources over factual recall. This paper introduces a novel evaluation framework that systematically assesses RALMs under three user need cases—Context-Exclusive, Context-First, and Memory-First—across three distinct context settings: Context Matching, Knowledge Conflict, and Information Irrelevant. By varying both user instructions and the nature of retrieved information, our approach captures the complexities of real-world applications where models must adapt to diverse user requirements. Through extensive experiments on multiple QA datasets, including HotpotQA, DisentQA, and our synthetic URAQ dataset, we find that restricting memory usage improves robustness in adversarial retrieval conditions but decreases peak performance with ideal retrieval results and model family dominates behavioral differences. Our findings highlight the necessity of user-centric evaluations in the development of retrieval-augmented systems and provide insights into optimizing model performance across varied retrieval contexts, explicitly separating factual correctness from faithfulness-to-instruction so readers know which dimension each score reflects. We will release our code and URAQ dataset upon acceptance of the paper.
@@ 1 Introduction
Recent advances in Language Models (LMs) have yielded impressive performance in knowledge-intensive tasks through Retrieval Augmented Generation (RAG) Lewis et al. (2020), including Real-time Question Answering Wang et al. (2024b), Educational Tutoring Han et al. (2024), and Personal Assistants Wang et al. (2024c). While these applications showcase RAG’s versatility, they also demand LMs that can adapt to diverse user needs—expressed via instructions on whether to prioritize external evidence or internal knowledge, or adhere strictly to specified (even potentially counter-factual) contexts, as seen in compliance, creative writing, or hypothetical scenarios. For instance, Real-time QA may rely heavily on updated external facts, whereas tutoring may draw more on the model’s conceptual understanding. Despite this potential, current RAG methods still struggle with identifying relevant references Laban et al. (2024), resolving knowledge conflicts Wang et al. (2024a), and reasoning effectively Islam et al. (2024). These challenges underscore the need for robust evaluation strategies capturing how well Retrieval Augmented Language Models (RALMs) adapt to evolving user requirements.
@@
Even though existing RAG/RALM benchmarks (Yu et al., 2024; Es et al., 2023; Chen et al., 2024)—including those that focus on multi-scenario evaluations (Friel et al., 2024; Zhu et al., 2024)—have advanced retrieval-augmented evaluation, they typically assume a single “optimal” approach to external information (e.g., always relying on retrieved context). This narrow perspective overlooks how diverse user instructions can dramatically alter the desired model’s behavior within the same scenario. In medical fact-checking, for instance, one user might demand answers derived only from peer-reviewed studies, while another relies on the model’s internal knowledge—even if these sources conflict (Miao et al., 2024). Such constraints underscore an urgent question: how can we systematically evaluate LMs under varying context usage requirements to reflect different user needs?
In this paper, we present a simple yet effective evaluation framework that rigorously examines how Retrieval-Augmented Language Models (RALMs) respond and be faithful to varying user instructions and context conditions. We consider three generic user cases(1) Context-Exclusive , (2) Context-First, and (3) Memory-First —to capture different degrees of reliance on external information versus internal knowledge. Alongside these cases, we vary the context settings(a) Context Matching, (b) Knowledge Conflict, and (c) Information Irrelevant—to represent scenarios where retrieved materials may align with, contradict, or fail to address the query. By intersecting user cases with distinct context conditions, we more closely mirror the complexities of real-world applications, where both the user’s priorities and the reliability of retrieved information can shift dramatically. This approach reveals how each scenario might alter the correct response—especially when context and memory conflict—an aspect often overlooked in previous work.
@@ 2.3 Recent RAG Benchmark
Previous RAG benchmarks like RAGAS Es et al. (2023) and RGB Chen et al. (2024) have facilitated progress by quantifying performance across various scenarios. However, many of these benchmarks focused on a single type of optimal setting in terms of context usages (for instance, always prioritizing the context), overlooking how different user instructions may drastically affect model behaviors and performances. Moreover, previous multi-scenario evaluations (Friel et al. 2024; Zhu et al. 2024), while covering a wide range of specific tasks and purpose abundant metrics for evaluating different aspects of RAG systems, also tend to follow the paradigm of focusing on singular optimality, neglecting that different user needs can actually happen in the same scenario, ultimately hindering the comprehensiveness of benchmark. FaithEval Ming et al. (2025) proposes a benchmark to evaluate the faithfulness of the RAG system. Our work diverges by decoupling evaluation criteria from predefined singular optimality and measuring model capability to adapt to dynamic user needs by using different instructions. This mirrors real-world deployments where systems must honor diverse users’ requirements rather than optimize for monolithic accuracy.
@@ 3.1 User Need Cases
Figure 2: An illustration of the framework with an example question with its possible retrieved context and the ground truth answer under each situation. According to different user needs and context settings, the ground truth answer can be different, reflecting instructed faithfulness (e.g., to ’Proxima Centauri’ if dictated by context and user need) rather than absolute factual correctness.
@@ 3.2 Context Settings
All retrieved contexts are unrelated to the question. This setting simulates the Needle-In-a-Haystack Laban et al. (2024) situation and tests the model’s ability on knowledge selection.
@@ 4.1 Datasets
Overview of QA Datasets
This experiment employs three QA datasets: HotpotQA Yang et al. (2018), DisentQA Neeman et al. (2023), and our synthetic User-focused Retrieval-Augmented QA (URAQ). To assess RALMs’ real-world performance, we use HotpotQA and DisentQA versions augmented with conflicted knowledge by Shaier et al. (2024) for the retrieval-content knowledge conflict setting. While valuable, these benchmarks lack controlled knowledge boundaries and have varying question difficulty, limiting evaluation. They also rely on long-document contexts only, thereby restricting retrieval diversity in terms of document length. In addition, the nature of factual-based for these datasets makes them may not be fully aligned with the evaluation of under needs. URAQ complements these by providing uniformly difficult questions and numerous concise modified contexts, specifically to isolate instruction-following and conflict-resolution capabilities when adapting to varied user needs, distinct from general comprehension over long or highly complex factual texts. While specialized domain datasets (e.g., medical, real-time QA) would be ideal for demonstrating our three user needs, we opted for these known and synthetic benchmarks to ensure reproducibility, broader comparability, and generalizability within budget constraints. The framework itself remains applicable to more domain-specific evaluations.
@@
Table 1: Basic information of the three datasets used in the experiment. The number of retrieved context is increased in a exponential way until the average number of tokens at the highest number of each sequence reaches around 20k in order to balance the effectiveness of the experiment on long context and the consumption of computational resources. The number of maximum tokens among all samples for a dataset may vary based on context retrieved.
URAQ Construction
We construct URAQ by first generating simple, distinct knowledge statements via GPT-4o-mini OpenAI et al. (2024) and removing near-duplicates using SentenceBERT Reimers and Gurevych (2019), then creating both original and “manipulated” versions by substituting key information or adding negations. For each knowledge pair, we produce a question requiring 1–5 reasoning steps and two separate answers (one from the original knowledge, one from the manipulated), ultimately selecting the 4-hop subset for the final dataset. A detailed description of this procedure is provided in Appendix A. To ensure fairness in evaluating multiple models, which may possess different internal knowledge, our experiments (particularly with URAQ) utilize a subset of questions for which the underlying correct factual knowledge is confirmed to be known by all evaluated models. This is achieved by pre-screening models on single-hop versions of questions related to the original knowledge. This pipeline ensures applicability across various domains and enables users to convert any datasets that previously designed for factuality and truthfulness into a faithfulness-oriented dataset, which adapts to our experiment on user needs.
4.2 Retrieval Context Setup
To examine how performance changes with varying amounts of retrieved context, rather than using a fixed retrieval count as in previous work Zhu et al. (2024), we evaluate LM performance by exponentially increasing the retrieval count across different datasets, shown in Table 1. To assess the models’ tolerance to distracting or irrelevant contexts, we ensure that only one relevant context is present for both the context-matching and conflicting settings, randomly positioned within the prompt. A detailed description of the prompt formatting and example is in the Appendix C. All other contexts are selected from a pool of original and manipulated knowledge that excludes any information directly related to the current question.
@@ 4.3 Evaluation Metrics
To rigorously assess user-need awareness across different user needs with different retrieval content, we test each user need with identical questions but varying the guidance on context usage, spanning three levels:
@@ 5.2 General Performance for Each User Need
(1) Case-Level Accuracy curve of Qwen2.5 on HotpotQA.
(2) Case-Level Accuracy curve of Llama-3.1 on HotpotQA.
Figure 4: Case-Level Accuracy curve of Qwen2.5 and Llama-3.1 on HotpotQA.
@@ 5.3.1 Performance on Optimal Context
Table 2: Percentage of errors that is "I don’t know" among the shortest 100 randomly selected samples that under Context Matching setting that is incorrect for Context-Exclusive user need and correct for Context or Memory-First.
@@ 6 Conclusion
We introduce an evaluation framework for RALMs that systematically assesses performance across diverse user needs and context settings. By decomposing user instructions into three generic user need cases (Context-Exclusive, Context-First, Memory-First) and three context settings (Context Matching, Knowledge Conflict, Information Irrelevant), our framework provides comprehensive insights into model capabilities and limitations. Our analysis covers overall user requirements, case-level evaluations, and the impact of varying context contents across different context lengths. The findings highlight the need for user-centric evaluations and architectural innovations to enhance RAG system reliability and real-world applicability.
@@ B.3 Memory-First
Appendix C Input Prompt Formatting
[…]
where <sys></sys> and <user></user> denote the system prompt and the user prompt. Among all data samples, the $I_{u}$ and $C$ may change according to the user case and context setting, while the $I_{f}$ remaining the same by instructing models to directly output a simple answer that is either a numeric value, a boolean ("yes" or "no"), or an entity, as described in Section 4.
We introduce an example input prompt that is designed for Context-Exclusive and Context Matching with 2 total retrieved context following the abstract input $(I_{f},I_{u},C,Q)$. The prompt is formatted with XML for both input and output. Specifically, the formatting instructions $I_{f}$ are separated into two parts: 1) The first and second instructions in the system prompt describing that the answer should be as simple as possible with XML format. 2) The instruction in the user prompt about format of context with an reinforcement of output format. The user need instruction $I_{u}$ is at the third instruction in the system prompt. The retrieved context $C$ is all the sentences in the user prompt within <info> xml tag. Finally, the question $Q$ is at the last of user prompt within <question> xml tag.